OpenStudy (anonymous):

Simplify x squared minus 36 all over x squared minus 12 times x plus 36..

OpenStudy (anonymous):

can,t understand ur question till now?

OpenStudy (anonymous):

(x^2 - 36) / (x^2 - 12x + 36) = (x - 6)(x + 6) / (x - 6)^2 = (x + 6) / (x - 6) a^2 - b^2 = (a + b)(a - b) a^2 + 2ab + b^2 = (a + b)^2